OpenStudy (anonymous):

does anyone know how to work this problem out? it says solve this equation by the method of substitution: 2x-4y=-14 -x-6y=-5

mathslover (mathslover):

\[\Huge{2x-4y=-14}\] First solve this equation for x \[\Huge{2x=-14+4y}\] \[\Huge{x=\frac{-14-4y}{2}}\] \[\Huge{x=-7-2y}\]

mathslover (mathslover):

now put the value of x that you get in the second equation ... can you do it now ?

OpenStudy (anonymous):

let me try it!

mathslover (mathslover):

ok

OpenStudy (anonymous):

for the x= -7-2y do i divide by -2 to cancel the 2y out?

mathslover (mathslover):

no why r u cancelling .. let it be there

mathslover (mathslover):

\[\Huge{x-6y=-5}\] \[\Huge{(-7-2y-6y=-5}\]

mathslover (mathslover):

solve this now

OpenStudy (anonymous):

oh! sorry! im so confused :(

OpenStudy (anonymous):

how do i find out the value of x, so i can plug it in the second equation?

mathslover (mathslover):

Ok let me start a new solution for you ..

mathslover (mathslover):

just see the first equation

mathslover (mathslover):

Can you solve for x ?

OpenStudy (anonymous):

don't you do -2x?

mathslover (mathslover):

right subtract 2x from both sides what you get ?

OpenStudy (anonymous):

-4y= 16?

mathslover (mathslover):

see if you subtract 2x on LHS then you should also on RHS \[\Huge{2x-2x-4y=-14-2x}\] \[\Huge{-4y=-14-2x}\]

mathslover (mathslover):

Now solve for x :: can you do it now ?

OpenStudy (anonymous):

i got -4y+2x=-14

OpenStudy (anonymous):

is probably not right

mathslover (mathslover):

\[\Huge{-4y+14=-2x}\]

OpenStudy (anonymous):

oh ok

mathslover (mathslover):

then divide both sides by (-2)

OpenStudy (anonymous):

i mean i know the answer but i just want someone teach me step by step on how to get that answer

mathslover (mathslover):

did u get this yet..

OpenStudy (anonymous):

divide both sides which are?

OpenStudy (anonymous):

-2x and -4y?

mathslover (mathslover):

-4y+14 and -2x

mathslover (mathslover):